OpenStudy (anonymous):

The mesures of the interior angles of a triangle are represented by 2x+10 and 3x+40. Which of the following is the mesure of the smallest angle of this triangle? 1.)50 2.)88 3.)42 4.)16

OpenStudy (hba):

But there should be 3 sides ?

OpenStudy (anonymous):

oh oops my bad its 2x+10, 5x-30 and 3x+40.

OpenStudy (hba):

Do you know what is the sum of angles ?

OpenStudy (anonymous):

no I dont know any of this... I dont get it

OpenStudy (hba):

Sum of interior angles=(n-2)*180 Where n=3 Now find Sum of interior angles :)

OpenStudy (anonymous):

what?

OpenStudy (hba):

Sum of interior angles=(3-2)*180 =

OpenStudy (hba):

What will Sum of interior angles be ?

OpenStudy (anonymous):

idk

OpenStudy (hba):

No one know until you do the maths here -_- Sum of interior angles=(3-2)*180 =1*180 2x+10+ 5x-30 + 3x+40 =180 Now solve for x.

OpenStudy (anonymous):

okay so then whats x?

OpenStudy (hba):

Do it your self,no one is gonna tell.

OpenStudy (anonymous):

is it 50?

OpenStudy (hba):

No

OpenStudy (anonymous):

20?

OpenStudy (hba):

Stop fooling around and solve it

OpenStudy (anonymous):

x=20 right?

OpenStudy (hba):

I am not telling.

OpenStudy (hba):

Solve it and show me the working.

OpenStudy (anonymous):

you add all the xs then divide

OpenStudy (hba):

Yes add them and show me what you got.

OpenStudy (anonymous):

i got 16

OpenStudy (anonymous):

is that right?

OpenStudy (hba):

Yes :)

OpenStudy (anonymous):

thanks(:

OpenStudy (anonymous):

16 is not the answer

OpenStudy (anonymous):

oh then how do i get the answer?

OpenStudy (anonymous):

16 is the value of x, you still need to plug it into your angle equations.

OpenStudy (anonymous):

Plug 16 into the x of: 2x+10, 5x-30 and 3x+40. See which is smallest.

OpenStudy (anonymous):

so it would be 2(16+20) and would you multiply?

OpenStudy (anonymous):

\[ 2(16)+10 = ? \\ 5(16)-30 = ? \\ 3(16)+40 = ? \]

OpenStudy (anonymous):

Also you should give hba a medal.

OpenStudy (anonymous):

42 50 88

OpenStudy (anonymous):

so it would be 42?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

thanks